# The unrealized potential of agroforestry for an emissions-intensive agricultural commodity

### v1

@@ Abstract
Reconciling agricultural production with climate-change mitigation and adaptation is one of the most formidable problems in sustainability. One proposed strategy for addressing this problem is the judicious retention of trees in agricultural systems. However, the magnitude of the current and future-potential benefit that trees contribute remains uncertain, particularly in the agricultural sector where trees can also limit production. Here we help to resolve these issues across a West African region responsible for producing $\approx$60of the world’s cocoa, a crop that contributes one of the highest per unit carbon footprints of all foods. We use machine learning to generate spatially-explicit estimates of shade-tree cover and carbon stocks across the region. We find that existing shade-tree cover is low, and not spatially aligned with climate threat. But we also find enormous unrealized potential for the sector to counterbalance a large proportion of their high carbon footprint annually, without threatening production. Our methods can be applied to other globally significant commodities that can be grown in agroforests, and align with accounting requirements of carbon markets, and emerging legislative requirements for sustainability reporting.
@@ Introduction
Trees in agricultural systems are predicted to be one of the most effective nature-based solutions for climate change mitigation (through carbon sequestration) and adaptation (through climate buffering), and are also associated with higher levels of biodiversity and the provision of other ecosystem services [16, 40, 14, 80]. Nevertheless, the magnitude of the current benefit that trees bring for climate and conservation, as well as their potential future benefit, has been controversial and remains uncertain [61, 10, 80, 76, 87, 13]. This uncertainty persists even as 40% of non-Annex I countries (typically low- and middle- income countries) have included agroforestry as a key component of their Nationally Determined Contributions (NDCs) to reducing greenhouse gas emissions under the Paris Agreement[69], and even as numerous large-scale and industry-specific initiatives aim to increase the number and cover of trees both locally and globally [19, 42]. There is, therefore, an urgent need to develop cost-effective, accurate, and repeatable methods for generating high-resolution assessments of the status of trees in agricultural systems [80]. Accounting requirements for carbon markets, as well as emerging sustainability reporting requirements for corporations, only make this need more urgent [33, 9, 27].
@@
Here we generate high-resolution, spatially-explicit maps of shade-tree cover and biomass (carbon storage) across 7.04 million hectares of West African agricultural systems currently responsible for producing $\approx$60 % of the global supply of cocoa [35]. Quantifying the current and future status of trees in agricultural systems is particularly important given their sheer scale in the historically carbon-dense and biodiverse tropics and subtropics, and the unique potential of agroforestry for allowing climate-smart, sustainable agricultural intensification for a wide range of crops [16, 45]. Cocoa production systems themselves are an important test case. On the one hand, cocoa production has been responsible for widespread historical and contemporary deforestation, including in protected areas [45, 67, 12], and cocoa contributes to one of the most emissions-intensive footprints of all foods [68, 62]. On the other hand, studies done on small scales demonstrate that cocoa can be produced together with moderate densities of shade trees (up to  40 % cover), with substantial benefits for carbon sequestration and biodiversity, and crucially, with little or no loss of yield [16, 4, 26, 65, 15]. In combination, these impacts and features of cocoa production systems have led to a number of ambitious industry commitments based around the planting of millions of shade-trees in agroforests to improve the sustainability of the sector [24, 89, 90]. Yet, while the nature of cocoa production is central to the socio-economic, climate, and biodiversity futures of this entire region, the current status and future potential of these systems is unknown.
@@
To solve this problem we use machine learning to generate estimates of the cover and biomass of trees in cocoa agroforestry systems across Ghana and Côte d’Ivoire in West Africa. We generate maps of shade-tree cover and biomass using information in publicly-available satellite imagery, recent global maps of tree-canopy height [50], and an extensive ground-truth field campaign conducted over two years across the two countries (Methods). Our novel approach generates high-resolution, spatially-explicit estimates, tailored to a single crop growing across two countries and five different agro-climatic zones, and includes estimates of mapping uncertainty. This information enables the first assessment of the status of agroforestry adoption in cocoa production systems. Then, by generating an empirical relationship between shade-tree cover and biomass, we also generate the first estimates of the current status and future potential of carbon sequestration in cocoa production systems across the region. Our results demonstrate the enormous and currently unrealized potential for climate-smart, sustainable intensification for this important global commodity. Moreover, beyond consequences for cocoa production, we have developed methods and a workflow to assess the status of trees in agricultural systems that is repeatable and generally applicable across the tropics and subtropics, including for other globally significant commodities that can be productively and sustainably grown in agroforests (e.g. coffee, tea, rubber, oil palm, etc.).
@@
Levels of shade-tree cover are low, and misaligned with spatially-variable climate threats
Despite the widely acknowledged benefits of agroforestry for cocoa production and long-standing industry commitments to increase shade tree cover in these systems, we found that cocoa production is overwhelmingly dominated by full-sun monocultures and low-shade agroforestry (Fig. 1). Average shade-tree cover across both countries was $13.2\pm 9.2\%$ (mean±$\pm$±standard deviation; Ghana: $13.8\pm 9.2\%$; Côte d’Ivoire: $12.8\pm 9.2\%$). More than 80% of cocoa production is implemented on farms with less than 20% shade, and only 5.6% (396,422 ha) of cocoa is grown under shade levels at or above 30% cover (Fig. 1). Farms with the highest levels of shade tree cover (above 30 %) tend to occur in small, isolated patches, particularly in the east of Ghana and the north-west of Côte d’Ivoire (Fig. 1), and do not tend to overlap with recently deforested areas [45]. In general, observed cover levels do not compare favourably with common recommendations of 20 %, 30 %, and even 40 % cover for cocoa production systems in this region [72].
Surprisingly, levels of shade-tree cover are not only low, but also do not vary substantially across larger spatial scales, including between administrative regions and agro-climatic zones. For example, average cover levels across agro-climatic zones range between 11.5-15% (Fig. S1), representing a maximum average difference of only 3.5% between the zones. The highest average cover occurred in the moderate-very dry zone (Type 1) in the northern cocoa growing area, with cover levels gradually declining through the dry and wet zones, reaching the lowest average cover in the very hot-very wet coastal zone (Type 5). Specifically, shade cover decreased from 15.0% in the moderate-very dry zone to 13.6%, 12.4%, and 11.5% in progressively wetter zones, while the cold, wet zone (Type 4) had an average cover of 14.1%. Similar small differences were found between administrative regions with average cover levels ranging between 9 and 17% (Fig. S3). The lack of substantial differences across larger spatial scales is surprising for two reasons. First, large differences in the timing of the introduction of cocoa production across regions might have been expected to result in differences in shade cover. For example, agricultural transformation occurred later in the west than the east of Ghana, which might have been expected to result in lower levels of average cover in cocoa growing regions in the west of the country [47, 81]. Second, higher levels of average cover might have been expected to occur in areas that experience hotter drier climates (Type 1 and 2 climates), which are also the areas where the persistence of the crop is most vulnerable to climate change (Fig. S1)[73]. Higher average temperatures are predicted to limit cocoa production in the future, emphasizing the importance of the ability of shade-trees to buffer the effects of climate on the crop [73, 16, 60]. The absence of higher shade-tree cover in these more vulnerable areas suggests that farmers are not currently capitalising on the climate buffering benefits of shade trees, nor are they maintaining higher shade tree densities as a source of alternative income in these areas[74, 73, 21].
Despite the low levels of shade-tree cover that we demonstrate, some of our estimates compare favourably with those used in official reporting. The Land Cover Map of Ghana [48] provides the only other estimates of shade-tree cover across this region, but this map resolves land use only coarsely and distinguishes only between unshaded (< 15 % cover) and shaded (>15 % cover) cocoa systems in line with Ghana’s REDD+ requirements (Ashiagbor, personal communication[30, 6]). Based on this map and classification, 13.6% of the cocoa growing area was considered "shaded" in 2021. By contrast, our map suggests that 37.2% of the cocoa-growing region in Ghana meets the REDD+ requirement of 15% cover, a nearly 3-fold increase over the existing estimate. A discrepancy of this magnitude is not surprising given our technical advance, and while our results demonstrate that shade-tree cover is generally low, this large positive discrepancy may nevertheless have important policy and economic implications.
@@
Figure 1: Map of shade-tree cover across cocoa-growing areas in Côte d’Ivoire and Ghana for the year 2022. The underlying data, estimated from Sentinel-2 imagery, has a 10 m ground sampling distance, but the map presented here was resampled to 100 m × 100 m for visualization purposes. The map of shade-tree cover is masked to cocoa-growing areas, as estimated by Kalischek et al. [45]. The histogram illustrates the distribution of shade-tree cover based on the original 10 m × 10 m resolution data. Location A highlights an area of cocoa monoculture in the Western Region, while Location B depicts an area of cocoa agroforestry in the Eastern Region of Ghana based on the original 10 m × 10 m resolution map.
@@
Cocoa production systems have enormous potential for climate-change mitigation
@@
One of the major benefits of trees in agricultural systems is climate change mitigation through carbon sequestration [25, 92, 58, 80, 40]. Past research on small scales has shown that cocoa agroforests can sequester substantial amounts of carbon in the aboveground biomass of trees, and that these benefits increase with increasing shade tree cover [58, 16, 60, 7]. Here we extend these small-scale studies to estimate carbon storage across the sector in this region. We find that cocoa production systems across Côte d’Ivoire and Ghana currently store an estimated $117.7\pm 1.2\times 10$-5 million tonnes of carbon (estimate $\pm$ 95 % CI) in above ground biomass. The majority of this carbon is distributed across a large area with relatively low carbon density, averaging $13.7\pm 6.9$ tonnes C ha-1 (estimate $\pm$ SD; Figs. 2,  S4).
@@
The role of forests in the sustainability of production landscapes
@@
Estimates of carbon storage across cocoa growing areas are high relative to unforested areas (Fig. 2), partly because cocoa itself is a medium-sized tree that also contributes to carbon sequestration . But because average shade-tree cover is low (Fig. 1), estimates of existing carbon densities and even future potential carbon sequestration capacity in these systems pale in comparison to those in the small remaining patches of intact forests in the region (Figs. 2S4). Indeed, although the area under cocoa production is more than three and a half times larger than the area of remaining intact forests, it stores only $\approx$40more carbon (Fig. 2, Fig. S4). This is because of large differences in carbon density between these two systems (intact forest: mean carbon density = $36.6\pm 19.7$ tonnes C $ha^{-}1$; cocoa: mean carbon density = $13.7\pm 6.9$ tonnes C $ha^{-}1$, Fig. S4). Even if shade-tree cover levels were increased to 30% across the entire cocoa growing region, the additional carbon 89 million tonnes C (Fig. 3) would represent ’only’ a little more than what is currently stored in the much smaller area of remaining intact forests (82.7 million tonnes C, Fig. S4c). These numbers emphasize the critical importance of intact forests (or ’avoided forest conversion’[40]) for climate change mitigation, even before accounting for their contribution to biodiversity and other ecosystem services [88]. In this context, recent findings about the continuing contribution of cocoa production to deforestation, including in protected areas, are all the more concerning [45, 67, 12].
@@
Figure 2: Map of aboveground biomass density (AGBD) across Côte d’Ivoire and Ghana for the year 2022. The underlying data product, estimated from Sentinel-2 imagery, has a 50-meter ground sampling distance. Location A illustrates AGBD patterns in a cocoa monoculture in the Western Region of Ghana. Location B depicts an area of cocoa agroforestry in the Eastern Region, while Location C shows undisturbed forest in Kakum National Park in the Central Region of Ghana. White areas indicate pixels for which no input image was available, often due to persistent cloud cover.
@@
Figure 3: Current aboveground carbon stocks and climate mitigation potential of cocoa systems across Ghana and Côte d’Ivoire. Bars represent aboveground carbon stocks under current conditions and projected scenarios with increased shade-tree cover (15%, 20%, 30%, and 40%). Numbers above the bars indicate the additional carbon stored compared to current levels. Due to the large number of estimates (over $3\times 10^{7}$ from our maps), the confidence intervals for current and projected values are too narrow to be visualized effectively. The 95% Confidence Interval for current carbon stocks is ±$\pm 1.2\times 10$-5. For the future projections, the 95% Highest Density Intervals (HDIs) are ±$\pm 2.0\times 10$-2, ±$\pm 4.2\times 10$-2, ±$\pm 8.3\times 10$-2, and ±$\pm 1.2\times 10$-1 million tonnes C for 15%, 20%, 30%, and 40% shade-tree cover adoption scenarios, respectively. The HDIs represent the range within which the means are statistically likely to reside.
@@
Our findings highlight the enormous potential for climate-change mitigation offered by the widespread implementation of cocoa agroforestry across Ghana and Côte d’Ivoire. While not measured directly in our study, known positive relationships between shade-tree cover and climate adaptation [16, 53], and shade-tree cover and biodiversity [16, 14, 15], suggest significant additional sustainability benefits would also emerge, with limited or no loss of cocoa yield [16, 4, 65, 15, 5]. However, the effort required to achieve these positive outcomes should not be underestimated. The Cocoa and Forest Initiative (CFI), which is one of the most successful industry initiatives for combating deforestation and promoting sustainable cocoa production, resulted in 31.1 million shade trees planted across approximately 570,000 hectares (assuming planting density 54.6 trees ha-1) between 2018 and 2022 [91, 89, 90]. Despite this effort, this amounts to just 8.1% of the cocoa growing area in this region [45]. This does suggest that, in the short-term, such initiatives should be targeted to areas where maximum benefits can be achieved, such as regions particularly vulnerable to climate threats, or where other sustainability benefits such as supporting biodiversity can be most effectively achieved. More generally, the adoption of agroforestry by small-holder farmers faces several barriers. These include the lack of immediately clear and obvious societal and economic incentives for including trees in production systems, as well as the availability of currently forested fertile land that can be rapidly converted into new production areas [45, 70]. Moreover, any carbon-sequestration benefits of including more trees in production systems will also be subject to concerns about permanence [40, 3], including in relation to insecure land tenure, and economic, governance, and climate risks [70, 83, 66, 3]. Clearly, realizing the large estimated benefits we describe will require an equally large effort in implementation. Regardless of the nature and success of this effort, however, the methods we have developed provide a baseline and tools for monitoring changes in the effectiveness of agroforestry interventions over time.
@@
Our maps and methods offer several advantages compared to previous work. First, we use a state-of-the-art cocoa map [45] to mask the shade cover map, ensuring that misidentification of landuse does not cause areas of shade-tree cover in these systems to be over- or under-estimated. Second, our shade map predicts shade levels with low error ($\approx$5%, Fig. S5a), and provides continuous shade levels, allowing users to employ and compare different thresholds between "shaded" and "unshaded" cocoa. Beyond compatibility with different specifications, the continuous scale increases reporting transparency, which is particularly important in the light of the ongoing debate about the definition of cocoa agroforestry, and varying targets set by individual companies and certification bodies. Third, our biomass map has higher resolution ($50\times 50$ meters) than the widely used European Space Agency’s (ESA) Climate Change Initiative (CCI; $100\times 100$ meters). Moreover, we take particular care to avoid overestimating low above-ground biomass density (AGBD), which can often occur as a consequence of efforts to mitigate the underestimation of high biomass. By contrast, our approach minimizes these errors in cocoa-growing areas, which is particularly important given that cocoa production systems are predominantly characterized by low biomass (Supplementary methods, Figs.2S4b). Finally, we aim for our project to support the implementation of zero-deforestation and carbon sequestration initiatives, particularly those relying on remote monitoring. As the demand grows for benchmarking commercial solutions, especially in relation to smallholder inclusion and carbon accounting, we provide our data, code, and maps openly. This transparency allows stakeholders to use current maps, but also to generate updated maps in the future, offering a cost-effective way for monitoring the long-term success of agroforestry interventions [32], and for meeting the demands of carbon accounting and sustainability reporting[33].
@@
We note three limitations to our work, with implications for future research. First, our methods required us to distinguish between shade trees and cocoa trees based on differences in their height (Methods). While our eight-meter cutoff minimizes the risk of misclassifying cocoa as shade trees, it could conceivably lead to the underestimation of shade-tree cover by excluding shade trees with canopies below eight meters in height. We think this risk is low given available estimates of shade-tree height in the study region [17]. Moreover, counter to this concern, our estimates for the total area of cocoa growing under 15% is substantially larger than previous estimates (cf. Land Cover Map of Ghana, described above). It is important to note, however, that newly-planted shade trees typically require 5-10 years to reach a height of eight meters, and so our findings do not account for the potential influence of recent plantings on future shade-tree cover. Second, because our biomass estimates (AGBD) are based on the canopy height estimates of Lang et al. [50], we inherit the overly smooth appearance of that map. As a result, the effective resolution of our AGBD map may be lower than its $50\times 50$ meter ground sampling distance. We do not expect these effects to adversely affect our conclusions, but caution against quantitative analyses and interpretations at small scales. Finally, while we focus on shade-tree cover, our approach does not consider the full complexity of agroforestry systems, including species composition, structural diversity, or information on farm management, for example [78, 17]. We believe a focus on cover is justified because it is consistent with previous work and agroforestry recommendations, directly influences solar insolation to the crop, can be related to biomass, and can be estimated from satellite images as we have shown [16, 93]. However, using remote sensing to facilitate large-scale assessments of other important features of agroforests is an important area for future research.
Conclusions
Our work shows that low levels of shade-tree cover in cocoa agroforests across a region generating $\approx$60of global production currently represents a substantial missed opportunity for sustainable intensification for a sector often associated with large climate impacts and biodiversity loss [45]. This is particularly true given recent work showing production can be maintained at levels more than double the current average shade levels [16]. By overcoming barriers to implementation and more effectively meeting the industry’s own commitments, our results suggest that the cocoa sector may have a relatively unique opportunity to simultaneously meet ambitious climate, sustainability, and economic goals. Similar progress in sustainability may be possible for other globally significant commodities that can be grown in agroforests, where assessments similar to our own may be highly beneficial. Finally, our study also further emphasizes the importance of a focus on maximising production within existing producing areas, so that the disproportionately large climate and other benefits of intact forests can be maintained.

### v2

@@ Abstract
[…]
@@ Introduction
Integrating trees into agricultural systems has been widely recognized as one of the most effective nature-based solutions, offering co-benefits for climate-change mitigation through carbon storage, climate-change adaptation via microclimate regulation, and support for biodiversity and ecosystem services [blaser18, griscom2017, bennett2022, terasaki2023]. Nevertheless, the magnitude of the current benefit that trees bring for climate and conservation, as well as their potential future benefit, has been controversial and remains uncertain [phalan2011, balmford2018, terasaki2023, sills2019, veldman2019, bastin2019]. This uncertainty persists even though agroforestry is included in the Nationally Determined Contributions (NDCs) of around 40% of non-Annex I countries under the Paris Agreement, as a strategy to reduce greenhouse gas emissions[rosenstock2019], and despite numerous large-scale and industry-specific initiatives aiming to increase the number and cover of trees both locally and globally [brancalion2020, holl2020]. There is, therefore, an urgent need to develop cost-effective, accurate, and repeatable methods for generating high-resolution assessments of the status of trees in agricultural systems [terasaki2023]. Accounting requirements for carbon markets, as well as emerging sustainability reporting requirements for corporations (e.g. EU, UK, and Australia), only make this need more urgent.
@@
Here we generate high-resolution, spatially-explicit maps of shade-tree cover and biomass (carbon storage) across 7.04 million hectares of West African agricultural systems currently responsible for producing $\approx$60% of the global supply of cocoa [FAO22]. Quantifying the current and future status of trees in agricultural systems is particularly important given their sheer scale in the historically carbon-dense and biodiverse tropics and subtropics, and the unique potential of agroforestry for allowing climate-smart, sustainable agricultural intensification for a wide range of crops [blaser18, kalischek23natfood]. Cocoa production systems themselves are an important test case. On the one hand, cocoa production has been responsible for widespread historical and contemporary deforestation, including in protected areas [kalischek23natfood, Renier_2023, barima2016], and cocoa contributes to one of the most emissions-intensive footprints of all foods (Ritchie (2020)[poore2018]. On the other hand, studies done on small scales demonstrate that cocoa can be produced together with moderate densities of shade trees (up to $\approx$30-50% cover), with substantial benefits for carbon sequestration and biodiversity, and crucially, with little or no loss of yield [blaser18, andres2018, clough2011, ramirez-argueta2022, bisseleua2009, schroth2016c]. In combination, these impacts and features of cocoa production systems have led to a number of ambitious industry commitments based around the planting of millions of shade-trees in agroforests to improve the sustainability of the sector [carodenuto2021](e.g. Cocoa & Forests Initiative [CFI]). Yet, while the nature of cocoa production is central to the socio-economic, climate, and biodiversity futures of this entire region, the current status and future potential of these systems is unknown.
@@
To solve this problem we use machine learning to generate estimates of the cover and biomass of trees in cocoa agroforestry systems across Ghana and Côte d’Ivoire in West Africa (Methods). We generate maps of shade-tree cover and biomass using information in publicly-available satellite imagery, recent global maps of tree-canopy height [lang2023high], and an extensive ground-truth field campaign conducted over two years across the two countries. Our approach generates high-resolution, spatially-explicit estimates, tailored to a single crop growing across two countries and five different agro-climatic zones, and includes estimates of mapping uncertainty. This information enables an assessment of the status of agroforestry adoption in cocoa production systems. Then, by generating an empirical relationship between shade-tree cover and biomass, we also generate estimates of the current status and future potential of carbon sequestration in cocoa production systems across the region (Methods). Our results demonstrate the enormous and currently unrealized potential for climate-smart, sustainable intensification for this important global commodity. Moreover, beyond consequences for cocoa production, we have developed methods and a workflow to assess the status of trees in agricultural systems that is repeatable and generally applicable across the tropics and subtropics, including for other globally significant commodities that can be productively and sustainably grown in agroforests (e.g. coffee).
@@
Shade-tree cover is low, and misaligned with climate threats
[…]
@@
Figure 1: Map of shade-tree cover across cocoa-growing areas in Côte d’Ivoire and Ghana for the year 2022. The shade-cover estimates were derived from Sentinel-2 imagery at a ground sampling distance of $10\times 10$ m. For visualization, the map presented here was resampled to a coarser grid of $100\times 100$ m. Predictions were restricted to areas mapped as cocoa-growing regions using the cocoa mask from Kalischek et al. (2023)[kalischek23natfood]. The histogram illustrates the distribution of shade-tree cover based on the original $10\times 10$ m resolution data ($n=679,593,915$ individual pixels) across all cocoa-growing areas. Location A depicts an area of cocoa monoculture in the Western Region, while Location B depicts an area of cocoa agroforestry in the Eastern Region of Ghana based on the original $10\times 10$ m resolution map. Base map boundaries were derived from the Global Administrative Areas database (GADM v4.1, https://gadm.org), used under academic license.
Surprisingly, and counter to expectations and previous research[knudsen2015, thompson2022, ruf2015], levels of shade-tree cover are not only low, but also do not vary substantially across larger spatial scales, including between administrative regions and agro-climatic zones. For example, average cover levels across agro-climatic zones range between 11.5-15% (Extended Data Fig. Extended Data 1), representing a maximum average difference of only 3.5% between the zones. The highest average cover occurred in the moderate-very dry zone (Type 1) in the northern cocoa growing area, with cover levels gradually declining through the dry and wet zones, reaching the lowest average cover in the very hot-very wet coastal zone (Type 5). Specifically, shade cover decreased from 15.0% in the moderate-very dry zone to 13.6%, 12.4%, and 11.5% in progressively wetter zones, while the cooler, wet zone (Type 4) had an average cover of 14.1%. Similar small differences were found between administrative regions with average cover levels ranging between 9 and 17% (Extended Data Fig. Extended Data 2). The lack of substantial differences across larger spatial scales is surprising for two reasons. First, differences in the timing of the introduction of cocoa production among regions might have been expected to result in differences in shade cover among regions[knudsen2015, thompson2022, ruf2015]. By contrast, the relative uniformity in cover that we document suggests that any such systematic differences across large scales have disappeared. This could be because of the widespread adoption of cocoa monocultures across all regions, and/or because of more spatially-targeted efforts to increase shade levels in areas most dominated by monocultures. Both processes could have reduced historical differences in cover between regions. Regardless of the mechanism, our results emphasize that low shade-tree cover is a pervasive issue that is not limited to recently established farms but affects cocoa-growing landscapes across both countries. Second, higher levels of average cover might have been expected to occur in areas that experience hotter climates as long as they are not too dry[Abdulai2018], which are also the areas where the persistence of the crop is most vulnerable to climate change (Extended Data Fig. Extended Data 1)[schroth2016b, schroth2017]. Higher average temperatures are predicted to limit cocoa production in the future, emphasizing the importance of the ability of shade-trees to buffer the effects of climate on the crop [schroth2016b, blaser18, niether2020]. The absence of higher shade-tree cover in these more vulnerable areas suggests that farmers are not currently capitalizing on the climate buffering benefits of shade trees, nor are they maintaining higher shade tree densities as a source of alternative income in these areas[schroth2017, schroth2016b, bunn19].
Despite the low levels of shade-tree cover that we demonstrate, some of our estimates compare favourably with those used in official reporting. The national land use map for Ghana produced by the Ghana Forestry Commission provides the only other estimates of shade-tree cover across this region, but this map resolves land use only coarsely and distinguishes only between unshaded (< 15% cover) and shaded (>15% cover) cocoa systems in line with Ghana’s REDD+ requirements (Ashiagbor, personal communication[epa2007, ashiagbor2020]). Based on this map and classification, 13.6% of the cocoa growing area was considered "shaded" in 2021. By contrast, our map suggests that 37.2% of the cocoa-growing region in Ghana meets the REDD+ requirement of 15% cover, a nearly 3-fold increase over the existing estimate. A discrepancy of this magnitude is not surprising given our technical advance, and while our results demonstrate that shade-tree cover is generally low, this large positive discrepancy may nevertheless have important policy and economic implications.
@@
Cocoa systems have major climate-change mitigation potential
@@
One of the major benefits of trees in agricultural systems is climate-change mitigation through carbon sequestration [chapman2020, zomer2016, nair2012, terasaki2023, griscom2017]. Past research on small scales has shown that cocoa agroforests can sequester substantial amounts of carbon in the aboveground biomass of trees, and that these benefits increase with increasing shade-tree cover [nair2012, blaser18, niether2020, asigbaase2021]. Here we extend these small-scale studies to estimate carbon storage across the sector in this region (Methods). We find that cocoa production systems across Côte d’Ivoire and Ghana currently store an estimated $117.7\pm 1.2\times 10$-5 million tonnes of carbon (estimate $\pm$ 95% CI) in above ground biomass. The majority of this carbon is distributed across a large area with relatively low carbon density, averaging $13.7\pm 6.9$ tonnes C ha-1 (estimate $\pm$ SD; Fig. 2, Extended Data Fig. Extended Data 3).
@@
Forests remain key to carbon storage in cocoa landscapes
@@
Estimates of carbon storage across cocoa growing areas are high relative to unforested areas (Fig. 2), partly because cocoa itself is a medium-sized tree that also contributes to carbon sequestration. But because average shade-tree cover in cocoa growing areas is low (Fig. 1), estimates of existing carbon densities and even future potential carbon sequestration capacity in these systems pale in comparison to those in the small remaining patches of intact forests in the region (Fig. 2, Extended Data Fig. Extended Data 3). Indeed, although the area under cocoa production is more than three and a half times larger than the area of remaining intact forests, it stores only $\approx$40% more carbon (Fig. 2, Extended Data Fig. Extended Data 3). This is because of large differences in carbon density between these two systems (mean $\pm$ standard deviation; intact forest: $36.6\pm 19.7$ tonnes C $ha^{-1}$; cocoa: $13.7\pm 6.9$ tonnes C $ha^{-1}$, Extended Data Fig. Extended Data 3). Even if shade-tree cover levels were increased to 30% across the entire cocoa growing region, the additional 84 million tonnes carbon (Fig. 3) would represent only a little more than what is currently stored in the much smaller area of remaining intact forests (82.7 million tonnes carbon, Extended Data Fig. Extended Data 3c). These numbers emphasize the critical importance of intact forests (or "avoided forest conversion" [griscom2017]) for climate-change mitigation, even before accounting for their contribution to biodiversity and other ecosystem services [watson2018]. In this context, findings about the continuing contribution of cocoa production to deforestation, including in protected areas, are all the more concerning [kalischek23natfood, Renier_2023, barima2016, ruf2015].
@@
Figure 2: Map of aboveground biomass density (AGBD) across Côte d’Ivoire and Ghana for the year 2022. The AGBD estimates were derived from Sentinel-2 imagery at a ground sampling distance of $50\times 50$ m. Location A illustrates AGBD patterns in a cocoa monoculture in the Western Region of Ghana. Location B depicts an area of cocoa agroforestry in the Eastern Region, while Location C shows undisturbed forest in Kakum National Park in the Central Region of Ghana. White areas indicate pixels for which no input image was available, often due to persistent cloud cover. Base map boundaries were derived from the Global Administrative Areas database (GADM v4.1, https://gadm.org), used under academic license.
@@
Figure 3: Current aboveground carbon stocks and climate mitigation potential of cocoa systems across Ghana and Côte d’Ivoire. Bars represent aboveground carbon stocks under current conditions and projected scenarios with increased shade-tree cover (15%, 20%, 30%, and 40%). Numbers above the bars indicate the additional carbon stored compared to current levels. Due to the large number of estimates (over $3\times 10^{7}$ from our maps), the confidence intervals for current and projected values are too narrow to be visualized effectively. The 95% Confidence Interval for current carbon stocks is $\pm 1.2\times 10$-5. For the future projections, the 95% Highest Density Intervals (HDIs) are $\pm 4.4\times 10$-2, $\pm 7.5\times 10$-2, $\pm 3.0\times 10$-1, and $\pm 7.0\times 10$-1 million tonnes C for 15%, 20%, 30%, and 40% shade-tree cover adoption scenarios, respectively. The HDIs represent the range within which the means are statistically likely to reside.
@@
Our findings highlight the enormous potential for climate-change mitigation offered by the widespread implementation of cocoa agroforestry across Ghana and Côte d’Ivoire. While not measured directly in our study, known positive relationships between shade-tree cover and climate adaptation [blaser18, lin2007], and shade-tree cover and biodiversity [blaser18, bennett2022, bisseleua2009], suggest significant additional sustainability benefits would also emerge, with limited or no loss of cocoa yield [blaser18, andres2018, ramirez-argueta2022, bisseleua2009, asare2018, schroth2016c]. However, the effort required to achieve these positive outcomes should not be underestimated. The Cocoa and Forest Initiative (CFI), which is one of the most successful industry initiatives for combating deforestation and promoting sustainable cocoa production, resulted in 31.1 million shade trees planted across approximately 570,000 hectares (assuming planting density 54.6 trees ha-1) between 2018 and 2022 (CFI progress report 2023, CFI initial action plans for Ghana and Côte d’Ivoire). Despite this effort, this amounts to just 8.1% of the cocoa growing area in this region [kalischek23natfood], and (assuming 30-40% cover at maturity) would result in an increase average shade-tree cover across the region from 13.2% to approximately 14.6-15.4%. While not insignificant (Fig. 3), this suggests that, in the short-term, such initiatives should be targeted to areas where maximum benefits can be achieved, such as regions particularly vulnerable to climate threats, or where other sustainability benefits such as supporting biodiversity can be most effectively achieved. More generally, the adoption of agroforestry by small-holder farmers faces several barriers, including limited short-term incentives, insecure land tenure, and the availability of forested fertile land that can be rapidly converted into new production areas [kalischek23natfood, ruf2011]. Moreover, the climate benefits of tree planting hinge on permanence [griscom2017, anderegg2020], which can be undermined by economic pressures, governance failures, or climate extremes [ruf2011, unruh2008, reichstein2013, anderegg2020]. Clearly, realizing the large estimated benefits we describe will require an equally substantial and sustained effort in implementation. This effort must prioritize maximizing production within existing cultivated areas, in order to avoid further loss of intact forests that offer disproportionately high climate and biodiversity value. However, regardless of the nature and success of this effort, the maps and methods developed here provide a baseline and practical tools for monitoring agroforestry adoption and carbon sequestration over time, supporting sustainability reporting and guiding targeted interventions.
@@
Our maps and methods offer several advantages compared to previous work. First, we use a state-of-the-art cocoa map[kalischek23natfood] to mask the shade-cover map, ensuring that misidentification of land use does not cause areas of shade-tree cover in these systems to be over- or under-estimated. Second, our shade map predicts shade levels with low error ($\approx$5%, Supplementary Fig. S1a), and provides continuous shade levels, allowing users to employ and compare different thresholds between ’shaded’ and ’unshaded’ cocoa. Beyond compatibility with different specifications, the continuous scale increases reporting transparency, which is particularly important in the light of the ongoing debate about the definition of cocoa agroforestry, and varying targets set by individual companies and certification bodies. Third, our biomass map has higher resolution ($50\times 50$ meters) than the widely used European Space Agency’s (ESA) Climate Change Initiative (CCI) map ($100\times 100$ meters; Supplementary Methods). Moreover, we take particular care to avoid overestimating low above-ground biomass density (AGBD), which can often occur as a consequence of efforts to mitigate the underestimation of high biomass. By contrast, our approach minimizes these errors in cocoa-growing areas (Supplementary Fig. S1b), which is particularly important given that cocoa production systems are predominantly characterized by low biomass (Fig. 2, Extended Data Fig. Extended Data 3b). Finally, we aim for our project to support the implementation of zero-deforestation and carbon sequestration initiatives, particularly those relying on remote monitoring. As the demand grows for benchmarking commercial solutions, especially in relation to smallholder inclusion and carbon accounting, we provide our data, code, and maps openly. This transparency allows stakeholders to use current maps, but also to generate updated maps in the future, offering a cost-effective way for monitoring the long-term success of agroforestry interventions, and for meeting the demands of carbon accounting and sustainability reporting.
@@
We note three limitations to our work, with implications for future research. First, our methods required us to distinguish between shade trees and cocoa trees based on differences in their height (Methods). While our eight-meter cutoff minimizes the risk of misclassifying cocoa as shade trees, it could conceivably lead to the underestimation of shade-tree cover by excluding shade trees with canopies below eight meters in height. We think this risk is low given available estimates of shade-tree height in the study region [blaser_2021]. Moreover, counter to this concern, our estimates for the total area of cocoa grown under 15% shade cover is substantially larger than previous estimates (cf. land use map for Ghana, described above). It is important to note, however, that newly-planted shade trees typically require 5-10 years to reach a height of eight meters, and so our findings do not account for the potential influence of recent plantings on future shade-tree cover. Second, because our biomass estimates (AGBD) are based on the canopy height estimates of Lang et al. [lang2023high], we inherit the overly smooth appearance of that map. As a result, the effective resolution of our AGBD map may be lower than its $50\times 50$ meter ground sampling distance. We do not expect these effects to adversely affect our conclusions, but caution against quantitative analyses and interpretations at small scales. Finally, while we focus on shade-tree cover, our approach does not consider the full complexity of agroforestry systems, including species composition, structural diversity, or information on farm management, for example [somarriba2024, blaser_2021]. We believe a focus on cover is justified because it is consistent with previous work and agroforestry recommendations, directly influences solar insolation to the crop, can be related to biomass, and can be estimated from satellite images as we have shown [blaser18, zuidema2005]. However, using remote sensing to facilitate large-scale assessments of other important features of agroforests is an important area for future research. More generally, extending our methods, and applying them to other globally significant commodities that can be grown in agroforests may be highly beneficial.